OpenStudy (anonymous):

Carmen is planning to invest $200 in a retirement account at the end of each month for the next 20 years. The account is earning 3.15% interest, compounded annually. He used the following formula and variables to solve for the future value of the account after 20 years. FVOA = Future Value of an Ordinary Annuity C = 200 n = 1 t = 20 i = 0.0315 He found that the future value of this account will be $5456.83. Is Carmen’s solution correct? If not, explain what he did wrong and provide the correct solution.

OpenStudy (anonymous):

Well n should aactually be 240 since it is 20yrs at 12 month's per year

OpenStudy (anonymous):

Payment should be -200 since it is a cash outflow

OpenStudy (anonymous):

the future value??

OpenStudy (anonymous):

You can use the geometric summation of: \[\sum_{n = 1}^{240} $200(1 + \frac{0.0315}{12})^n\]\[$200(1 + \frac{0.0315}{12})[\frac{1 - (1 + \frac{0.0315}{12})^{240}}{1 - (1 + \frac{0.0315}{12})}]\] Which comes out to approximately $66,922.75

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

Just use financial concept it's way easier

OpenStudy (anonymous):

ok i'll remember that

OpenStudy (anonymous):

I derived the summation from the compound interest formula \[A = P(1 + \frac{r}{n})^{nt}\] There is no other way to do the problem except with Calculus, geometric summation is the most simplest method

OpenStudy (anonymous):

oh ok got it

OpenStudy (anonymous):

Use ur financial calculator it's way faster

OpenStudy (anonymous):

i am now

OpenStudy (anonymous):

Did you get it

OpenStudy (anonymous):

dont know yet i'll find ouy in a couple of minutes